# Littlewood-Offord bounds on the symmetric groups and applications

### v1

@@ 2.4. Paper Organization
$\sum_{r\in{\mathcal{R}}}\Big\|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{C}n.$
@@
$\sum_{i,j,r}\Big\|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\gg_{\delta,C}\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}\leq\frac{1}{Cn}}\frac{t^{2}n^{4}D_{k}^{2}}{\Delta^{2}}|{\mathcal{G}}_{k}|+\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}>\frac{1}{Cn}}n|{\mathcal{G}}_{k}|.$
@@
Case 1: $\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}\leq\frac{1}{Cn}}D_{k}^{2}|{\mathcal{G}}_{k}|\asymp_{{\varepsilon}}n$.
@@
$\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}\leq\frac{1}{Cn}}\frac{t^{2}n^{4}D_{k}^{2}}{\Delta^{2}}|{\mathcal{G}}_{k}|\gg_{{\varepsilon}}An^{2}\log n.$
Case 2: $\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}>\frac{1}{Cn}}D_{k}^{2}|{\mathcal{G}}_{k}|\asymp_{{\varepsilon}}n$.
As $D^{2}_{k}\leq 1/(A^{2}\log n)$, we obtain $\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}>\frac{1}{Cn}}|{\mathcal{G}}_{k}|\gg_{{\varepsilon}}A^{2}n\log n$, which implies
$\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}>\frac{1}{Cn}}n|{\mathcal{G}}_{k}|\gg_{{\varepsilon}}A^{2}n^{2}\log n.$
@@
$\sum_{i,j,r}\|tn^{1/2}(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\gg_{{\varepsilon},\delta,C}An^{2}\log n.$
@@
$\frac{c_{\delta}}{n^{2}}\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{G}}_{0}\\ r\in{\mathcal{R}}\end{subarray}}\Big\|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{\delta}\frac{1}{n^{2}}\cdot|{\mathcal{G}}_{0}|\cdot n\asymp_{\delta,{\varepsilon}}n.$
@@
$\frac{c_{\delta}}{n^{2}}\sum_{i,j,r}\Big\|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\frac{c_{\delta}}{n^{2}}\sum_{i,j,r}\Big|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big|^{2}\asymp_{\delta,\Delta}t^{2}n^{3}.$
@@
Consider the case where $w_{1}=\dots=w_{(1-t)n}=0$ and $w_{(1-t)n+1}=\dots=w_{n}=\pm 1/\sqrt{\delta n}$, which satisfies the condition of Theorem 1.15. In this setting, it can be shown that the weighted sum $\sqrt{\delta}\sum_{i}w_{i}\pi(i)$ spreads rather evenly over the interval $Ct^{1/2}n^{3/2},Ct^{1/2}n^{3/2}]$ for some constant C>0C>0.666This follows because the variance of the sum is of order $tn^{3}$. By pigehole principle, the point anti-concentration of this integer-valued sum is $\Theta(t^{-1/2}n^{-3/2})$, $\sup_{x}{\mathbb{P}}(|\sqrt{tn}\sum_{i}w_{i}\pi(i)=x)\gg t^{-1/2}n^{-3/2})$. Hence for any $\delta>0$
$\sup_{x}{\mathbb{P}}(|\sum_{i}w_{i}\pi(i)/n-x|\leq\delta)\gg t^{-1/2}n^{-3/2}.$
@@
$\frac{c_{\delta}}{n^{2}}\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{G}}_{0}\\ r\in{\mathcal{R}}\end{subarray}}\Big\|tn^{1/2-{\varepsilon}}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{\delta}\frac{1}{n^{2}}\cdot|{\mathcal{G}}_{0}|\cdot n\asymp_{\delta,{\varepsilon}}n.$
@@
To conclude this section, we can further refine the scaling under an additional—but still quite generic—condition on the $w_{i}$, as follows.
@@
Let $0<{\varepsilon}<1/2$ and $\delta>0$ be given. Suppose that the sequence $(w_{1},\ldots,w_{n})$ satisfies Condition 1.14 for some sufficiently large constant $A>0$. In addition, assume that no interval of length ${\varepsilon}/\sqrt{n}$ contains more than $(1-{\varepsilon})n$ of the values $w_{i}$, and that there are at least $An\log n$ pairs $(i,j)$ for which
@@
${\mathbb{P}}\Big(\Big|\sum_{i}w_{i}\,\frac{\pi(i)}{n}-L\Big|\leq\frac{1}{n^{5/2-{\varepsilon}}}\Big)=O\left(\frac{1}{n^{5/2-{\varepsilon}}}\right).$
We can also obtain an $L$-dependent bound, as in Theorems 1.15 and 1.17. This result can be seen as a continuous analog of Theorem 1.24. Here, roughly speaking, if the $w_{i}$ are spread out evenly over the interval $C/\sqrt{n},\,C/\sqrt{n}]$, then the average consecutive spacing is $1/n^{3/2}$. The above condition requires that most of the consecutive spacings asymptotically attain this bound.
@@
$\sup_{L}{\mathbb{P}}(|n\sum_{i}w_{i}v_{\pi(i)}-Ln|\leq 1)\ll\int_{|t|\leq 1}|\varphi(t)|\,dt\ll\int_{|t|\leq 1}e^{-\frac{1}{n^{3}}\sum_{i,j,k,l}\|tn(w_{i}-w_{j})(v_{k}-v_{l})\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}}.$
@@
    $\sum_{r\in I}\|tw(P_{d}(r)-P_{d}(r_{0}))/n^{d-1}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{C}n.$
@@
$\frac{1}{n^{3}}\sum_{(i,j)\in{\mathcal{G}}_{k}}\sum_{r,r_{0}\in I}\|t(w_{i}-w_{j})(P_{d}(r)-P_{d}(r_{0}))/n^{d-1}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{C}\frac{1}{n}|{\mathcal{G}}_{k}|.$
@@
$\asymp\frac{1}{n^{3}}t^{2}D_{k}^{2}|{\mathcal{G}}_{k}|n^{4}\asymp t^{2}nD_{k}^{2}|{\mathcal{G}}_{k}|.$
@@
$\frac{1}{n^{3}}\sum_{(i,j)\in{\mathcal{G}}}\sum_{r,r_{0}\in I}\|t(w_{i}-w_{j})(P_{d}(r)-P_{d}(r_{0}))/n\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp\sum_{k:\,|t|D_{k}\leq 1/(Cn)}t^{2}nD_{k}^{2}|{\mathcal{G}}_{k}|+\frac{1}{n}\sum_{k:\,|t|D_{k}>1/(Cn)}|{\mathcal{G}}_{k}|.$
@@
$\asymp t^{2}n^{2},$
where we used the fact that $\sum_{1\leq i,j\leq n}(w_{i}-w_{j})^{2}=2n$, and that $\sum_{r,r_{0}\in I}\big(P_{d}(r)-P_{d}(r_{0})\big)^{2}\asymp n^{2d+2}$. We then have
@@
$\sum_{r\in I}\|t(w_{i}-w_{j})(P_{d}(r)-P_{d}(r_{0}))/n^{d-1}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp n.$
@@
$\sum_{1\leq i,j\leq n;\,r,r_{0}\in I}\|t(w_{i}-w_{j})(P_{d}(r)-P_{d}(r_{0}))/n^{d-1}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\gg n^{2}|{\mathcal{G}}|\gg A^{2}n^{3}\log n.$
@@
Let $b,B>0$ be constants. Suppose that the sequence $(w_{1},\ldots,w_{n})$ satisfies Condition 1.14 for some sufficiently large constant $A>0$. Let $I\subset[n]$ be any subset with $|I|\geq\delta n$, and consider the sequences $(v_{1},\dots,v_{n})$ and $(v^{\prime}_{1},\dots,v^{\prime}_{n})$ partially defined by
@@
By Fact 4.1, there exists a subset ${\mathcal{R}}\subset\{-n,\dots,n\}$ with $|{\mathcal{R}}|=\Theta_{\delta}(n)$, such that for every $h\in{\mathcal{R}}$, there are $\Theta_{\delta}(n)$ pairs $r,s\in I$ with rs=hr-s=h. For each $s\in I$, let $J_{s}$ denote the set of $h\in{\mathcal{R}}$ such that $s+h\in I$. By double counting the pairs $r,s\in I$ with $r-s\in{\mathcal{R}}$, we deduce that $|J_{s}|=\Theta_{\delta}(n)$ for $\Theta_{\delta}(n)$ many $s\in I$. Denote the set of such $s$ by $I_{0}$.
@@
$\sum_{r}\|t_{2}(w_{i}-w_{j})(r-s)\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{C}n,$
@@
$\sum_{r}\big\|t_{1}(w_{i}-w_{j})\frac{br^{2}+b_{n}^{\prime}r-bs^{2}-b_{n}^{\prime}s}{n}+t_{2}(w_{i}-w_{j})(r-s)\big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}\asymp_{C}n.$
@@
$\sum_{r\in I}|(t_{1},t_{2})\cdot(\frac{br^{2}+b_{n}^{\prime}r-bs^{2}-b_{n}^{\prime}s}{n^{2}},\frac{r-s}{n})|^{2}\asymp n.$
@@
$|\frac{b(h+2s)h+b_{n}^{\prime}h}{n^{2}}\frac{h^{\prime}}{n}-\frac{b(h^{\prime}+2s)h^{\prime}+b_{n}^{\prime}h^{\prime}}{n^{2}}\frac{h}{n}|=|\frac{b(h-h^{\prime})hh^{\prime}}{n^{3}}|\asymp 1.$
@@
$|(t_{1},t_{2})\cdot(\frac{b(h+2s)h+b_{n}^{\prime}h}{n^{2}},\frac{h}{n})|^{2}+|(t_{1},t_{2})\cdot(\frac{b(h^{\prime}+2s)h^{\prime}+b_{n}^{\prime}h^{\prime}}{n^{2}},\frac{h^{\prime}}{n})|^{2}\asymp 1.$
@@
$\sum_{r}(w_{i}-w_{j})^{2}|t_{1}\big(\frac{br^{2}+b_{n}^{\prime}r-bs^{2}-b_{n}^{\prime}s}{n^{2}}\big)+t_{2}(r-s)|^{2}\asymp(w_{i}-w_{j})^{2}(t_{1}^{2}+t_{2}^{2})n^{3}.$
@@
(of Lemma 6.5) If $1\leq|t_{1}|\leq\sqrt{A\log n}$, then we can argue as in the proof of (10) for very large $|t|$. If $(\sqrt{A\log n})/n\leq|t_{1}|\leq 1$, we can also argue as in the proof of (9) for large $|t|$. It remains to assume $|t_{1}|\leq(\sqrt{A\log n})/n$, in which case $1/2\leq|t_{2}|\leq\sqrt{A\log n}$. To this end, we recall (43) that $|t_{1}(w_{i}-w_{j})\frac{br^{2}+b_{n}^{\prime}r-bs^{2}-b_{n}^{\prime}s}{n}|\leq\frac{2}{\sqrt{A}}$. We can then argue as in the proof of (10), again applying Corollary 2.5. ∎
By Lemmas 6.2 and 6.5, it suffices to focus on $\|{\bm{t}}\|_{2}\leq\frac{\sqrt{A\log n}}{n}$. In this case,
@@
$\frac{1}{n^{2}}\sum_{r,s\in I}\big|t_{1}\frac{br^{2}+b_{n}^{\prime}r-bs^{2}-b_{n}^{\prime}s}{n}+t_{2}(r-s)\big|^{2}\asymp(t_{1}^{2}+t_{2}^{2})n^{2}.$
@@
${\mathbb{P}}(\|{\bm{X}}\|_{2}\leq 1)\leq\frac{1}{n^{2}}\int_{|x_{1}|\leq L_{2}}|f(x_{1})|dx_{1}=O(\frac{1}{n^{2}}L_{2}e^{-\Theta(L_{2}^{2})})=O(\frac{1}{n^{2}}e^{-\Theta(L_{2}^{2})}),$
@@
Our proof shows that we can actually relax the condition $|\sum_{i}v^{\prime}_{i}|\ll|\sum_{i}v_{i}|$ to $|\overline{{\bm{v}}^{\prime}}\cdot\overline{{\bm{w}}}|\ll|\overline{{\bm{v}}}\cdot\overline{{\bm{w}}}|$. In the centered case, where $\overline{{\bm{w}}}=0$, the latter condition holds for all $v_{i}$ and $v^{\prime}_{i}$. Note also that a weaker version of this result, in which the right-hand side is $O\Big(\frac{\sqrt{\log n}}{n}\Big)$, can be handled by a much simpler method (by not relying on Theorem 1.19 for the joint event, but instead using Theorem 1.18 together with Lemma 2.3). We leave the details to the reader.
@@
${\mathbb{P}}\big(|X_{2}|\in I_{k},\ |X_{1}|\in I_{l}/n\big)\ll n\cdot\frac{1}{n^{2}}e^{-\Theta((k-a)^{2})}e^{-\Theta((l/n-b)^{2})}\ll\frac{1}{n}e^{-\Theta((k-a)^{2})}e^{-\Theta((l/n-b)^{2})}.$
@@
${\mathbb{P}}\big(|X_{1}|\leq|X_{2}|/n\big)\ll\frac{1}{n}\sum_{1\leq|l|\leq|k|}e^{-\Theta((k-a)^{2})}e^{-\Theta((l/n-b)^{2})}.$
@@
$\frac{1}{n}\sum_{1\leq|l|\leq|k|<2|a|}e^{-\Theta((k-a)^{2})}e^{-\Theta((l/n-b)^{2})}\ll\frac{1}{n}\sum_{1\leq|l|\leq|k|<2|a|}e^{-\Theta(b^{2})}\ll\frac{a^{2}e^{-\Theta(b^{2})}}{n}\ll\frac{1}{n}$

### v2

@@ 4.3. Diophantine Properties
$\sum_{r\in{\mathcal{R}}}\Big\|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{C}(n).$
@@
$\sum_{i,j,r}\Big\|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Omega_{\delta,C}(1)\cdot\Big(\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}\leq\frac{1}{Cn}}\frac{t^{2}n^{4}D_{k}^{2}}{\Delta^{2}}|{\mathcal{G}}_{k}|+\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}>\frac{1}{Cn}}n|{\mathcal{G}}_{k}|\Big).$
@@
Case 1: $\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}\leq\frac{1}{Cn}}D_{k}^{2}|{\mathcal{G}}_{k}|=\Theta_{{\varepsilon}}(n)$.
@@
$\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}\leq\frac{1}{Cn}}\frac{t^{2}n^{4}D_{k}^{2}}{\Delta^{2}}|{\mathcal{G}}_{k}|=\Omega_{{\varepsilon}}(An^{2}\log n).$
Case 2: $\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}>\frac{1}{Cn}}D_{k}^{2}|{\mathcal{G}}_{k}|=\Theta_{{\varepsilon}}(n)$.
As $D^{2}_{k}\leq 1/(A^{2}\log n)$, we obtain $\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}>\frac{1}{Cn}}|{\mathcal{G}}_{k}|=\Omega_{{\varepsilon}}(A^{2}n\log n)$, which implies
$\sum_{k:\,\frac{|t|n^{1/2}}{\Delta}D_{k}>\frac{1}{Cn}}n|{\mathcal{G}}_{k}|=\Omega_{{\varepsilon}}(A^{2}n^{2}\log n).$
@@
$\sum_{i,j,r}\|tn^{1/2}(w_{i}-w_{j})r\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Omega_{{\varepsilon},\delta,C}(An^{2}\log n).$
@@
$\frac{c_{\delta}}{n^{2}}\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{G}}_{0}\\ r\in{\mathcal{R}}\end{subarray}}\Big\|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{\delta}\!\Big(\frac{|{\mathcal{G}}_{0}|}{n}\Big)=\Theta_{\delta,\varepsilon}(n).$
@@
$\frac{c_{\delta}}{n^{2}}\sum_{i,j,r}\Big\|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\frac{c_{\delta}}{n^{2}}\sum_{i,j,r}\Big|\frac{tn^{1/2}}{\Delta}(w_{i}-w_{j})r\Big|^{2}=\Theta_{\delta,\Delta}(t^{2}n^{3}).$
@@
Consider the case where $w_{1}=\dots=w_{(1-t)n}=0$ and $w_{(1-t)n+1}=\dots=w_{n}=\pm 1/\sqrt{\delta n}$, which satisfies the condition of Theorem 2.6. In this setting, one can show that the weighted sum $\sqrt{\delta n}\sum_{i}w_{i}\pi(i)$ spreads rather evenly over the interval $Ct^{1/2}n^{3/2},Ct^{1/2}n^{3/2}]$ for some constant C>0C>0.444This follows since the variance is of order $tn^{3}$. By the pigeonhole principle, $\sup_{x}{\mathbb{P}}(\sqrt{\delta n}\sum_{i}w_{i}\pi(i)=x)=\Omega(t^{-1/2}n^{-3/2})$. Hence for any $\delta>0$
$\sup_{x}{\mathbb{P}}(|\sum_{i}w_{i}\pi(i)/n-x|\leq\delta)=\Omega(t^{-1/2}n^{-3/2}).$
@@
$\frac{c_{\delta}}{n^{2}}\sum_{\begin{subarray}{c}(i,j)\in{\mathcal{G}}_{0}\\ r\in{\mathcal{R}}\end{subarray}}\Big\|tn^{1/2-{\varepsilon}}(w_{i}-w_{j})r\Big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{\delta}\Big(\frac{|{\mathcal{G}}_{0}|}{n}\Big)=\Theta_{\delta,{\varepsilon}}(n).$
@@
To conclude this section, we further refine the scaling under an additional—yet still quite generic—condition on the $w_{i}$, as stated below. The proof will be given in Appendix A.
@@
Let $0<{\varepsilon}<1/2$ and $\delta>0$ be given. Suppose that the sequence $(w_{1},\ldots,w_{n})$ satisfies Condition 2.5 for some sufficiently large constant $A>0$. In addition, assume that no interval of length ${\varepsilon}/\sqrt{n}$ contains more than $(1-{\varepsilon})n$ of the values $w_{i}$, and that there are at least $An\log n$ pairs $(i,j)$ for which
@@
${\mathbb{P}}\Big(\Big|\sum_{i}w_{i}\,\frac{\pi(i)}{n}-L\Big|\leq\frac{1}{n^{5/2-{\varepsilon}}}\Big)=O\Big(\frac{1}{n^{5/2-{\varepsilon}}}\Big).$
We can also obtain an $L$-dependent bound, as in Theorems 2.6 and 2.8. This result can be seen as a continuous analog of Theorem 3.3. Here, roughly speaking, if the $w_{i}$ are spread out evenly over the interval $C/\sqrt{n},\,C/\sqrt{n}]$, then the average consecutive spacing is $1/n^{3/2}$. The above condition requires that most of the consecutive spacings asymptotically attain this bound.
@@
$\sup_{L}{\mathbb{P}}(|n\sum_{i}w_{i}v_{\pi(i)}-Ln|\leq 1)=O\Big(\int_{|t|\leq 1}|\varphi(t)|\,dt\Big)=O\Big(\int_{|t|\leq 1}e^{-\frac{1}{n^{3}}\sum_{i,j,k,l}\|tn(w_{i}-w_{j})(v_{k}-v_{l})\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}}\Big).$
@@
    $\sum_{r\in I}\|tw(P_{d}(r)-P_{d}(r_{0}))/n^{d-1}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{C}(n).$
@@
$\frac{1}{n^{3}}\sum_{(i,j)\in{\mathcal{G}}_{k}}\sum_{r,r_{0}\in I}\|t(w_{i}-w_{j})(P_{d}(r)-P_{d}(r_{0}))/n^{d-1}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{C}\Big(\frac{|{\mathcal{G}}_{k}|}{n}\Big).$
@@
$\Theta\Big(\frac{1}{n^{3}}t^{2}D_{k}^{2}|{\mathcal{G}}_{k}|n^{4}\Big)=\Theta(t^{2}nD_{k}^{2}|{\mathcal{G}}_{k}|).$
@@
$\frac{1}{n^{3}}\sum_{(i,j)\in{\mathcal{G}}}\sum_{r,r_{0}\in I}\|t(w_{i}-w_{j})(P_{d}(r)-P_{d}(r_{0}))/n\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Omega(1)\cdot\Big(\sum_{k:\,|t|D_{k}\leq 1/(Cn)}t^{2}nD_{k}^{2}|{\mathcal{G}}_{k}|+\frac{1}{n}\sum_{k:\,|t|D_{k}>1/(Cn)}|{\mathcal{G}}_{k}|\Big).$
@@
$=\Theta(t^{2}n^{2}),$
where we used the fact that $\sum_{1\leq i,j\leq n}(w_{i}-w_{j})^{2}=2n$, and that $\sum_{r,r_{0}\in I}\big(P_{d}(r)-P_{d}(r_{0})\big)^{2}=\Theta(n^{2d+2})$. We then have
@@
$\sum_{r\in I}\|t(w_{i}-w_{j})(P_{d}(r)-P_{d}(r_{0}))/n^{d-1}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta(n).$
@@
$\sum_{1\leq i,j\leq n;\,r,r_{0}\in I}\|t(w_{i}-w_{j})(P_{d}(r)-P_{d}(r_{0}))/n^{d-1}\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Omega(n^{2}|{\mathcal{G}}|)=\Omega(A^{2}n^{3}\log n).$
@@
Let $b,B>0$ be constants. Suppose that the sequence $(w_{1},\ldots,w_{n})$ satisfies Condition 2.5 for some sufficiently large constant $A>0$. Let $I\subset[n]$ be any subset with $|I|\geq\delta n$, and consider the sequences $(v_{1},\dots,v_{n})$ and $(v^{\prime}_{1},\dots,v^{\prime}_{n})$ partially defined by
@@
By Fact 6.1, there exists a subset ${\mathcal{R}}\subset\{-n,\dots,n\}$ with $|{\mathcal{R}}|=\Theta_{\delta}(n)$, such that for every $h\in{\mathcal{R}}$, there are $\Theta_{\delta}(n)$ pairs $r,s\in I$ with rs=hr-s=h. For each $s\in I$, let $J_{s}$ denote the set of $h\in{\mathcal{R}}$ such that $s+h\in I$. By double counting the pairs $r,s\in I$ with $r-s\in{\mathcal{R}}$, we deduce that $|J_{s}|=\Theta_{\delta}(n)$ for $\Theta_{\delta}(n)$ many $s\in I$. Denote the set of such $s$ by $I_{0}$.
@@
$\sum_{r}\|t_{2}(w_{i}-w_{j})(r-s)\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{C}(n),$
@@
$\sum_{r}\big\|t_{1}(w_{i}-w_{j})\frac{br^{2}+b_{n}^{\prime}r-bs^{2}-b_{n}^{\prime}s}{n}+t_{2}(w_{i}-w_{j})(r-s)\big\|_{{\mathbb{R}}/{\mathbb{Z}}}^{2}=\Theta_{C}(n).$
@@
$\sum_{r\in I}|(t_{1},t_{2})\cdot(\frac{br^{2}+b_{n}^{\prime}r-bs^{2}-b_{n}^{\prime}s}{n^{2}},\frac{r-s}{n})|^{2}=\Theta(n).$
@@
$|\frac{b(h+2s)h+b_{n}^{\prime}h}{n^{2}}\frac{h^{\prime}}{n}-\frac{b(h^{\prime}+2s)h^{\prime}+b_{n}^{\prime}h^{\prime}}{n^{2}}\frac{h}{n}|=|\frac{b(h-h^{\prime})hh^{\prime}}{n^{3}}|=\Theta(1).$
@@
$|(t_{1},t_{2})\cdot(\frac{b(h+2s)h+b_{n}^{\prime}h}{n^{2}},\frac{h}{n})|^{2}+|(t_{1},t_{2})\cdot(\frac{b(h^{\prime}+2s)h^{\prime}+b_{n}^{\prime}h^{\prime}}{n^{2}},\frac{h^{\prime}}{n})|^{2}=\Theta(1).$
@@
$\sum_{r}(w_{i}-w_{j})^{2}|t_{1}\big(\frac{br^{2}+b_{n}^{\prime}r-bs^{2}-b_{n}^{\prime}s}{n^{2}}\big)+t_{2}(r-s)|^{2}=\Theta\Big((w_{i}-w_{j})^{2}(t_{1}^{2}+t_{2}^{2})n^{3}\Big).$
@@
(of Lemma 8.5) If $1\leq|t_{1}|\leq\sqrt{A\log n}$, then we can argue as in the proof of (10) for very large $|t|$. If $(\sqrt{A\log n})/n\leq|t_{1}|\leq 1$, we can also argue as in the proof of (9) for large $|t|$. It remains to assume $|t_{1}|\leq(\sqrt{A\log n})/n$, in which case $1/2\leq|t_{2}|\leq\sqrt{A\log n}$.
To this end, we recall (43) that $|t_{1}(w_{i}-w_{j})\frac{br^{2}+b_{n}^{\prime}r-bs^{2}-b_{n}^{\prime}s}{n}|\leq\frac{2}{\sqrt{A}}$. We can then argue as in the proof of (10), again applying Corollary 4.5. ∎
By Lemmas 8.2 and 8.5, it suffices to focus on $\|{\bm{t}}\|_{2}\leq\frac{\sqrt{A\log n}}{n}$. In this case,
@@
$\frac{1}{n^{2}}\sum_{r,s\in I}\big|t_{1}\frac{br^{2}+b_{n}^{\prime}r-bs^{2}-b_{n}^{\prime}s}{n}+t_{2}(r-s)\big|^{2}=\Theta\Big((t_{1}^{2}+t_{2}^{2})n^{2}\Big).$
@@
${\mathbb{P}}(\|{\bm{X}}\|_{2}\leq 1)\leq\frac{1}{n^{2}}\int_{|x_{1}|\leq L_{2}}|f(x_{1})|dx_{1}=O\Big(\frac{1}{n^{2}}L_{2}e^{-\Theta(L_{2}^{2})}\Big)=O\Big(\frac{1}{n^{2}}e^{-\Theta(L_{2}^{2})}\Big),$
@@
Our proof shows that we can actually relax the condition $|\sum_{i}v^{\prime}_{i}|=O(|\sum_{i}v_{i}|)$ to $|\overline{{\bm{v}}^{\prime}}\cdot\overline{{\bm{w}}}|=O(|\overline{{\bm{v}}}\cdot\overline{{\bm{w}}}|)$. In the centered case, where $\overline{{\bm{w}}}=0$, the latter condition holds for all $v_{i}$ and $v^{\prime}_{i}$. Note also that a weaker version of this result, in which the right-hand side is $O(\frac{\sqrt{\log n}}{n})$, can be handled by a much simpler method (by not relying on Theorem 2.10 for the joint event, but instead using Theorem 2.9 together with Lemma 4.3). We leave the details to the reader.
@@
${\mathbb{P}}\big(|X_{2}|\in I_{k},\,|X_{1}|\in I_{l}/n\big)=O\Big(\frac{1}{n}e^{-\Theta((k-a)^{2})}e^{-\Theta((l/n-b)^{2})}\Big).$
@@
${\mathbb{P}}\big(|X_{1}|\leq|X_{2}|/n\big)=\frac{O(1)}{n}\sum_{1\leq|l|\leq|k|}e^{-\Theta((k-a)^{2})}e^{-\Theta((l/n-b)^{2})}.$
@@
$\frac{1}{n}\sum_{1\leq|l|\leq|k|<2|a|}e^{-\Theta((k-a)^{2})}e^{-\Theta((l/n-b)^{2})}\leq\frac{1}{n}\sum_{1\leq|l|\leq|k|<2|a|}e^{-\Theta(b^{2})}\leq\frac{16a^{2}e^{-\Theta(b^{2})}}{n}=O\Big(\frac{1}{n}\Big)$